OpenStudy (anonymous):

What's equal to (sqrt7)^3?

OpenStudy (anonymous):

\[\huge (\sqrt 7)^3 = 7^{{\frac {1}{2}} \times 3}= 7^{{\frac {3}{2}} }\]

OpenStudy (anonymous):

Thsi is the index form of the given problem...

OpenStudy (anonymous):

thank youu!